# Serializing Java Objects in Plain Code

### v1

@@ Abstract
To address this problem, we propose plain-code serialization. Our core idea is to serialize objects observed at runtime in the native syntax of a programming language. We realize this vision in the context of Java, and demonstrate a prototype which serializes Java objects to Java source code. The resulting source faithfully reconstructs the objects seen at runtime. Our prototype is called ProDJ and is publicly available. We experiment with ProDJ to successfully plain-code serialize $174,699$ objects observed during the execution of 4 open-source Java applications. Our performance measurement shows that the performance impact is not noticeable.
@@ 1 Introduction
Listing 1: Serialized objects in XML are not meant for human consumption, they are not readable. In this paper, we aim to serialize objects in plain code instead, as demonstrated later in 8
@@ 3 COSDictionary receiver = deserializeFromXML(
4␣␣<org.apache.pdfbox.cos.COSDictionary>
5␣␣␣␣<direct>false</direct>
6␣␣␣␣<needToBeUpdated>false</needToBeUpdated>
7␣␣␣␣<itemsclass=\”org.apache.pdfbox.util.SmallMap\”>
8␣␣␣␣␣␣<mapArr>
9␣␣␣␣␣␣␣␣<org.apache.pdfbox.cos.COSName>
10␣␣␣␣␣␣␣␣␣␣<direct>false</direct>
11␣␣␣␣␣␣␣␣␣␣<name>BaseFont</name>
12␣␣␣␣␣␣␣␣␣␣<hashCode>-1657267488</hashCode>
13␣␣␣␣␣␣␣␣</org.apache.pdfbox.cos.COSName>
14␣␣␣␣␣␣␣␣<org.apache.pdfbox.cos.COSName>
15␣␣␣␣␣␣␣␣␣␣<direct>false</direct>
16␣␣␣␣␣␣␣␣␣␣<name>Times-Roman</name>
17␣␣␣␣␣␣␣␣␣␣<hashCode>1058166902</hashCode>
18␣␣␣␣␣␣␣␣</org.apache.pdfbox.cos.COSName>
19␣␣␣␣␣␣</mapArr>
20␣␣␣␣</items>
21␣␣</org.apache.pdfbox.cos.COSDictionary>
22␣␣);
@@ 23 COSName key = deserializeFromXML(
24␣␣<org.apache.pdfbox.cos.COSName>
25␣␣␣␣<direct>false</direct>
26␣␣␣␣<name>BaseFont</name>
27␣␣␣␣<hashCode>-1657267488</hashCode>
28␣␣</org.apache.pdfbox.cos.COSName>
29␣␣);
@@
34}
To solve the problem of unreadable serialized data, we propose the novel concept of plain-code serialization, which consists in serializing an object directly into code statements. By serializing objects to plain code, the serialized representation becomes readable, using a syntax that is both concise and natural to developers. In addition, the compiler gives us strong guarantees about the static correctness of the serialized objects. Deserialization then simply means executing the code. In the context of tests, this completely removes dirty data like XML strings, thus making the generated tests more readable and self-contained [17].
@@
We demonstrate the feasibility of plain-code serialization by evaluating ProDJ with $4$ well-known Java projects. ProDJ successfully serializes $174,699$ Java objects observed at runtime, producing Java code that can reconstruct them. We find that the objects are faithfully reconstructed through a full-cycle of plain-code serialization and deserialization.
@@ 2.1 Concept & Overview
Benefits. Plain code serialization has three major benefits: readability, performance, and correctness checking. First, the serialized objects are readable for developers, using the programming language syntax for which they have been trained for. Second, deserialization is fast with no overhead, since it does not involve any I/O or reflection, which is the basis for most mainstream serialization libraries. Last but not least, when serializing to a statically typed language, plain-code serialization provides guarantees about the serialized data. In this case, the compiler checks for the validity of the object with the full power of the type system, capturing domain-specific correctness that one typically does not have with binary or markup serialization.
@@
We now illustrate how these two strategies work and complement each other. Consider the Monkey and Habitat classes in 2. Monkey is serializable using the structure-based strategy. The three possible reconstruction plans for the Monkey class are presented in 3. To create the Java code for a given Monkey object, the \$age\$, \$eyeColor\$, and \$habitat\$ meta-variables are replaced with actual values from runtime. However, note that the Habitat class in 2 contains the private field area. This field is part of the internal state of Habitat and, according to the information hiding principle, not directly exposed. Therefore, we use trace-based object serialization for Habitat. For example, recording the constructor invocation of Habitat, as well as an invocation of its grow method would allow us to reconstruct a Habitat object with coordinates "42, 42" and an area of 2.0. In the following subsections, we present the structure-based and trace-based object reconstruction strategies in detail.
Listing 2: The example Monkey class to be serialized, and its associated types
@@ 2.2.1 Synthesizing Reconstruction Plans
An action within a reconstruction plan is a statement that sets zero or more fields within an object. The $5$ action types used for structure-based object reconstruction are denoted as sb in Table 1. Each action can be constructing ($\alpha$) or field-setting ($\beta_{f}$ sets a field $f$). Constructor calls and calls to factory methods are constructing actions, whereas field assignments and setter calls are field-setting actions. We analyze a target class to extract all possible actions that can be combined in different ways to produce reconstruction plans. Each action has one or more meta-variables. A meta-variable is a placeholder for a field that is being set through the action. When a plan is instantiated to serialize a runtime object, each meta-variable is replaced with the actual value of the field.
@@
Let us illustrate the generation of a reconstruction plan using the Monkey class in 2. A Monkey has an integer value for age, and an eyeColor, which takes possible values from an enum. The Monkey also has a Habitat. Each field of Monkey has a corresponding meta-variable, namely \$age\$, \$eyeColor\$, and \$habitat\$. There are three possible reconstruction plans for Monkey, which we present in 3. The most idiomatic way to create a Monkey is using the first plan, which contains a single constructing action. It creates a Monkey by calling the three-argument age, color, and habitat constructor. The second and third plans consist of one constructing and two field-setting actions. The second plan calls the age constructor, followed by the eyeColor setter, and finally directly assigns the habitat field. The last plan has actions to call the age constructor, and directly assign both the eyeColor and habitat fields.
Listing 3: Three possible reconstruction plans for Monkey
@@ 2.2.3 Instantiating Reconstruction Plans
Finally, the instantiated reconstruction plans are converted to a plain-code representation by iterating over the actions (Table 1) and converting them one-by-one. A “call constructor” action first plain-code serializes each argument of the constructor and then emits a constructor call statement. Similarly, the arguments to the “call factory method” and “call method” actions are plain-code serialized before emitting a method call statement. “Assign field” plain-code serializes the value of the field and emits a field assignment statement, while “use enum constant” emits a read to the matching constant. We illustrate this procedure in 4 by reconstructing a Monkey using the first reconstruction plan from 3, which is the most idiomatic. For our Monkey, this means emitting a code statement that calls the three-argument constructor. The first argument, age, is a primitive and is directly inserted. The other two arguments are recursively plain-code serialized, stored as variables and referenced in the generated constructor call. The “use enum constant” action is used when plain-code serializing the eyeColor of the Monkey. Lastly, the “use object reference” defers serialization to the next strategy, trace-based plain-code serialization, and is explained in subsection 2.3. The habitat field of the Monkey is handled by this action type.
Listing 4: Structure-based plain-code serialization of a Monkey from 2
@@ 2.3.1 Capturing Events
Given a type we want to serialize, we first identify all types needed for its reconstruction because they are used either as fields, constructor parameters, or method parameters. Then, we instrument the constructors, accessible methods, and field assignments for all these types. For example, in the Habitat class (2), we instrument the constructor on line $28$, as well as the grow method on lines $30-33$. We show the resulting instrumented code for the grow method in 5. When the instrumented grow is invoked at runtime, it emits multiple events, ordered by a monotonically increasing counter representing logical time.
Listing 5: Instrumentation for monitoring events within the trace-based object reconstruction strategy, the injected code is highlighted in green.
@@
The first event, MethodStartEvent, marks the start of a method call, and contains all information necessary to convert it to a method call action: the receiver object, the fully qualified name of the invoked method, as well as all arguments to the method call. As method invocations can be nested, and methods can be recursive, we also track when a method invocation ends through the emission of a MethodEndEvent when the method returns. The start and end events share a common unique ID, which is used to pair them up. We also assign unique identifiers to each object. This ensures that all method invocations, constructor calls, and field assignments for an object can be associated. The object ID is used whenever an object is referenced within an event. In addition to the object ID, every reference also stores the current logical time, as an object may have different states at different points in time. We also see on line $5$ of 5 that the grow method updates the value of the area field. This corresponds to a FieldSetEvent event, which is emitted whenever a field of an object is reassigned inside a method. Within the FieldSetEvent, we store the ID of the receiver object, as well as the new and old values of the field being updated. Using FieldSetEvents, we can determine whether a given method invocation mutated the receiver object, and would need to be replayed during object reconstruction. Besides capturing method invocations and field assignments, we also instrument all constructors. Whenever a constructor finishes executing, a ConstructEvent is emitted. This event contains all necessary information to create a constructor call action, including the fully qualified name of the invoked constructor, as well as all of its arguments. Additionally, the event also contains the ID of the newly created object, allowing later events to refer to it. Thus, exercising the instrumented program with a workload emits events. The resulting captured event sequence is persisted on disk.
@@ 2.3.2 Mapping Events to Actions
Whenever the plain-code serialization of an object is requested, we follow the same procedure as for structure-based serialization. We illustrate this by plain-code serializing an instance of the Habitat class from 2. The resulting code is shown in 6. First, the constructor call action is converted to code by plain-code serializing all constructor arguments. In our example, this means serializing the coordinate parameter of the constructor. Next, a constructor call is emitted and the newly created object is stored in a local variable. This is done on line $2$ of 6. Then, all method call actions up until the requested logical time are also converted to code by plain-code serializing the parameters and then calling the method on the created object. For our Habitat this is a single method call, grow(42), on line $4$.
Listing 6: Trace-based plain-code serialization of a Habitat from 2
@@ 2.3.3 Object Reconstruction Mixing
We show such a case in 7, where we are constructing a Monkey. Monkey is serialized using structure-based serialization, and its habitat field is serialized using trace-based serialization. As the captured events are mapped to code offline, but the plain-code serialized habitat field has to be inserted into a specific spot in the generated code, the structure-based serialization strategy uses the “use object reference” (from Table 1) marker action. This action emits an $id@logical-time$ reference, which is replaced by the trace-based serialized value after the event sequence conversion is finished. Besides this action, Table 1 also contains two more actions, “use static field” and “use standard charset”. These are actions that can not be done structurally, but also do not need the full trace-based serialization machinery. For the “use static field” action type, we look at all static fields of the class we want to serialize and detect if any of these store a reference-equal object. If such a field can be found, we emit a field read instead of tracing the constructor and subsequent mutating method calls. A similar, but more specific, optimization is realized by the “use standard charset” action type. This action type replaces references to Java’s standard charsets with the canonical named constants, and serves as an example of integrating domain logic into the serialization process. To further facilitate extensibility, such custom adapters and action types can be provided as plugins.
Listing 7: Plain code serialization of a Monkey, mixing structure- and trace-based serialization
@@ 2.4.4 Implementation
ProDJ relies on Spoon [19] for source code analysis, in order to identify the associated types of methods involving serialization within the target Java application. ProDJ uses Z3 [20] to solve the constraint problem of subsubsection 2.2.2 and determine the most idiomatic reconstruction plan for types that are structure-based serializable. The event sequences for trace-based serializable objects are logged in JSON. The instrumentation of MUTs and their associated types is achieved using ByteBuddy [^3] https://bytebuddy.net/.
@@ 3.2 Subjects
PDFBox 8f23f8791 1,390 738 183 188
GraphHopper f64c57c2e 945 590 199 114
BroadLeaf 951216e942 2,933 1,883 508 444
Gephi 82684f927 1,057 1,721 140 225
@@
Table 2: Projects used in the evaluation of plain code serialization: number of serialization points we consider (#SPs), the number of types related to the serialization points (#Types), the number of serialization triggered in production, by our workloads (#SPs Prod) and the number of types that we target for plain code serialization at runtime (#Types Prod)
@@
For each project in Table 2, column $2$ links to the exact version (SHA) we use for our evaluation. Columns $3$ and $4$ give the number of methods instrumented by ProDJ within the project (#SPs for ‘serialization point’), and consequently, the number of associated types (#Types) that need to be instrumented at this point. For example, ProDJ adds serialization points in $2,933$ methods within BroadLeaf, involving $2,223$ types. Columns $5$ and $6$ of Table 2 provide the same metrics, but computed after runtime observations, i.e., the number of serialization points and types that are covered at runtime, which we discuss in the next subsection.

### v3

@@ Abstract
To address this problem, we propose plain-code serialization. Our core idea is to serialize objects observed at runtime in the native syntax of a programming language. We realize this vision in the context of Java, and demonstrate a prototype which serializes Java objects to Java source code. The resulting source faithfully reconstructs the objects seen at runtime. Our prototype is called ProDJ and is publicly available. We experiment with ProDJ to successfully plain-code serialize $174,699$ objects observed during the execution of 4 open-source Java applications. Our performance measurement shows that the performance impact is not noticeable. Through a user study, we demonstrate that developers prefer plain-code serialized objects within automatically generated tests over their representations as XML or JSON.
@@ 1 Introduction
Listing 1: Serialized objects in XML are not meant for human consumption, they are not readable. In this paper, we aim to serialize objects in plain code instead, as demonstrated later in 2 and 9
1@Test
@@ 4 COSDictionary receiver = deserializeFromXML(
5␣␣<org.apache.pdfbox.cos.COSDictionary>
6␣␣␣␣<direct>false</direct>
7␣␣␣␣<needToBeUpdated>false</needToBeUpdated>
8␣␣␣␣<itemsclass=\”org.apache.pdfbox.util.SmallMap\”>
9␣␣␣␣␣␣<mapArr>
10␣␣␣␣␣␣␣␣<org.apache.pdfbox.cos.COSName>
11␣␣␣␣␣␣␣␣␣␣<direct>false</direct>
12␣␣␣␣␣␣␣␣␣␣<name>BaseFont</name>
13␣␣␣␣␣␣␣␣␣␣<hashCode>-1657267488</hashCode>
14␣␣␣␣␣␣␣␣</org.apache.pdfbox.cos.COSName>
15␣␣␣␣␣␣␣␣<org.apache.pdfbox.cos.COSName>
16␣␣␣␣␣␣␣␣␣␣<direct>false</direct>
17␣␣␣␣␣␣␣␣␣␣<name>Times-Roman</name>
18␣␣␣␣␣␣␣␣␣␣<hashCode>1058166902</hashCode>
19␣␣␣␣␣␣␣␣</org.apache.pdfbox.cos.COSName>
20␣␣␣␣␣␣</mapArr>
21␣␣␣␣</items>
22␣␣</org.apache.pdfbox.cos.COSDictionary>
23␣␣);
@@ 24 COSName key = deserializeFromXML(
25␣␣<org.apache.pdfbox.cos.COSName>
26␣␣␣␣<direct>false</direct>
27␣␣␣␣<name>BaseFont</name>
28␣␣␣␣<hashCode>-1657267488</hashCode>
29␣␣</org.apache.pdfbox.cos.COSName>
30␣␣);
@@
35}
Listing 2: Automatically generated test containing plain-code serialized runtime objects receiver and key. Compared to the XML object representations in 1, this test is arguably more readable and maintainable.
1@Test
2@DisplayName(getNameAsString(org.apache.pdfbox.cos.COSName)”)
[…]
4 // Arrange
[…]
6 COSName key = COSName.BASE_FONT;
7 // Act
8 String actual = receiver.getNameAsString(key);
9 // Assert
10 assertThat(actual).isEqualTo(Times-Roman”);
11}
12
13COSDictionary createCOSDictionary1() {
14 // Recreated from trace
[…]
16 COSName key = COSName.TYPE;
17 COSBase value = COSName.FONT;
18 receiver.setItem(key, value);
19 COSName key0 = COSName.SUBTYPE;
20 COSBase value0 = COSName.TYPE1;
21 receiver.setItem(key0, value0);
22 COSName key1 = COSName.BASE_FONT;
23 receiver.setName(key1, Times-Roman”);
24 COSName key2 = COSName.ENCODING;
25 COSBase value1 = COSName.WIN_ANSI_ENCODING;
26 receiver.setItem(key2, value1);
[…]
28}
To solve the problem of unreadable serialized data, we propose the novel concept of plain-code serialization, which consists in serializing an object directly into code statements. By serializing objects to plain code, the serialized representation becomes readable, using a syntax that is both concise and natural to developers. We illustrate this with the semantically equivalent test of 2, which contains the same serialized objects as 1 represented as plain code, instead of XML. In addition, the compiler gives us strong guarantees about the static correctness of the serialized objects. Deserialization then simply means executing the code. In the context of tests, this completely removes dirty data like XML strings, thus making the generated tests more readable and self-contained [17].
@@
We demonstrate the feasibility of plain-code serialization by evaluating ProDJ with $4$ well-known Java projects. ProDJ successfully serializes $174,699$ Java objects observed at runtime, producing Java code that can reconstruct them. We find that the objects are faithfully reconstructed through a full-cycle of plain-code serialization and deserialization. Through a user study with 17 developers, we also demonstrate that the use of plain-code serialized objects contributes to the readability of automatically generated tests, relative to XML- or JSON-serialized objects.
@@ 2.1 Concept & Overview
Benefits. Plain code serialization has two major benefits: readability and correctness checking. First, the serialized objects are readable for developers, using the programming language syntax for which they have been trained for. Second, when serializing to a statically typed language, plain-code serialization provides guarantees about the serialized data. In this case, the compiler checks for the validity of the object with the full power of the type system, capturing domain-specific correctness that one typically does not have with binary or markup serialization.
@@
We now illustrate how these two strategies work and complement each other. Consider the Monkey and Habitat classes in 3. Monkey is serializable using the structure-based strategy. The three possible reconstruction plans for the Monkey class are presented in 4. To create the Java code for a given Monkey object, the \$age\$, \$eyeColor\$, and \$habitat\$ meta-variables are replaced with actual values from runtime. However, note that the Habitat class in 3 contains the private field area. This field is part of the internal state of Habitat and, according to the information hiding principle, not directly exposed. Therefore, we use trace-based object serialization for Habitat. For example, recording the constructor invocation of Habitat, as well as an invocation of its grow method would allow us to reconstruct a Habitat object with coordinates "42, 42" and an area of 2.0. In the following subsections, we present the structure-based and trace-based object reconstruction strategies in detail.
Listing 3: The example Monkey class to be serialized, and its associated types
@@ 2.2.1 Synthesizing Reconstruction Plans
An action within a reconstruction plan is a statement that sets zero or more fields within an object. The $5$ action types used for structure-based object reconstruction are denoted as sb in Table 1. Each action can be constructing ($\alpha$) or field-setting ($\beta_{f}$ sets a field $f$). Constructor calls and calls to factory methods are constructing actions, whereas field assignments and setter calls are field-setting actions. We analyze a target class to extract all possible actions that can be combined in different ways to automatically synthesize reconstruction plans. Each action has one or more meta-variables. A meta-variable is a placeholder for a field that is being set through the action. When a plan is instantiated to serialize a runtime object, each meta-variable is replaced with the actual value of the field.
@@
Let us illustrate the generation of a reconstruction plan using the Monkey class in 3. A Monkey has an integer value for age, and an eyeColor, which takes possible values from an enum. The Monkey also has a Habitat. Each field of Monkey has a corresponding meta-variable, namely \$age\$, \$eyeColor\$, and \$habitat\$. There are three possible reconstruction plans for Monkey, which we present in 4. The most idiomatic way to create a Monkey is using the first plan, which contains a single constructing action. It creates a Monkey by calling the three-argument age, color, and habitat constructor. The second and third plans consist of one constructing and two field-setting actions. The second plan calls the age constructor, followed by the eyeColor setter, and finally directly assigns the habitat field. The last plan has actions to call the age constructor, and directly assign both the eyeColor and habitat fields.
Listing 4: Three possible reconstruction plans for Monkey
@@ 2.2.3 Instantiating Reconstruction Plans
Finally, the instantiated reconstruction plans are converted to a plain-code representation by iterating over the actions (Table 1) and converting them one-by-one. A “call constructor” action first plain-code serializes each argument of the constructor and then emits a constructor call statement. Similarly, the arguments to the “call factory method” and “call method” actions are plain-code serialized before emitting a method call statement. “Assign field” plain-code serializes the value of the field and emits a field assignment statement, while “use enum constant” emits a read to the matching constant. We illustrate this procedure in 5 by reconstructing a Monkey using the first reconstruction plan from 4, which is the most idiomatic. For our Monkey, this means emitting a code statement that calls the three-argument constructor. The first argument, age, is a primitive and is directly inserted. The other two arguments are recursively plain-code serialized, stored as variables and referenced in the generated constructor call. The “use enum constant” action is used when plain-code serializing the eyeColor of the Monkey. Lastly, the “use object reference” defers serialization to the next strategy, trace-based plain-code serialization, and is explained in subsection 2.3. The habitat field of the Monkey is handled by this action type.
Listing 5: Structure-based plain-code serialization of a Monkey from 3
@@ 2.3.1 Capturing Events
Given a type we want to serialize, we first identify all types needed for its reconstruction because they are used either as fields, constructor parameters, or method parameters. Then, we instrument the constructors, accessible methods, and field assignments for all these types. For example, in the Habitat class (3), we instrument the constructor on line $28$, as well as the grow method on lines $30-33$. We show the resulting instrumented code for the grow method in 6. When the instrumented grow is invoked at runtime, it emits multiple events, ordered by a monotonically increasing counter representing logical time.
Listing 6: Instrumentation for monitoring events within the trace-based object reconstruction strategy, the injected code is highlighted in green.
@@
The first event, MethodStartEvent, marks the start of a method call, and contains all information necessary to convert it to a method call action: the receiver object, the fully qualified name of the invoked method, as well as all arguments to the method call. As method invocations can be nested, and methods can be recursive, we also track when a method invocation ends through the emission of a MethodEndEvent when the method returns. The start and end events share a common unique ID, which is used to pair them up. We also assign unique identifiers to each object. This ensures that all method invocations, constructor calls, and field assignments for an object can be associated. The object ID is used whenever an object is referenced within an event. In addition to the object ID, every reference also stores the current logical time, as an object may have different states at different points in time. We also see on line $5$ of 6 that the grow method updates the value of the area field. This corresponds to a FieldSetEvent event, which is emitted whenever a field of an object is reassigned inside a method. We assume that the principle of encapsulation is followed, and fields are either all public for simple data classes, or not mutated from outside the class itself. Within the FieldSetEvent, we store the ID of the receiver object, as well as the new and old values of the field being updated. Using FieldSetEvents, we can determine whether a given method invocation mutated the receiver object, and would need to be replayed during object reconstruction. Besides capturing method invocations and field assignments, we also instrument all constructors. Whenever a constructor finishes executing, a ConstructEvent is emitted. This event contains all necessary information to create a constructor call action, including the fully qualified name of the invoked constructor, as well as all of its arguments. Additionally, the event also contains the ID of the newly created object, allowing later events to refer to it. Thus, exercising the instrumented program with a workload emits events. The resulting captured event sequence is persisted on disk.
@@ 2.3.2 Mapping Events to Actions
Whenever the plain-code serialization of an object is requested, we follow the same procedure as for structure-based serialization. We illustrate this by plain-code serializing an instance of the Habitat class from 3. The resulting code is shown in 7. First, the constructor call action is converted to code by plain-code serializing all constructor arguments. In our example, this means serializing the coordinate parameter of the constructor. Next, a constructor call is emitted and the newly created object is stored in a local variable. This is done on line $2$ of 7. Then, all method call actions up until the requested logical time are also converted to code by plain-code serializing the parameters and then calling the method on the created object. For our Habitat this is a single method call, grow(42), on line $4$.
Listing 7: Trace-based plain-code serialization of a Habitat from 3
@@ 2.3.3 Object Reconstruction Mixing
We show such a case in 8, where we are constructing a Monkey. Monkey is serialized using structure-based serialization, and its habitat field is serialized using trace-based serialization. As the captured events are mapped to code offline, but the plain-code serialized habitat field has to be inserted into a specific spot in the generated code, the structure-based serialization strategy uses the “use object reference” (from Table 1) marker action. This action emits an $id@logical-time$ reference, which is replaced by the trace-based serialized value after the event sequence conversion is finished. Besides this action, Table 1 also contains two more actions, “use static field” and “use standard charset”. These are actions that can not be done structurally, but also do not need the full trace-based serialization machinery. For the “use static field” action type, we look at all static fields of the class we want to serialize and detect if any of these store a reference-equal object. If such a field can be found, we emit a field read instead of tracing the constructor and subsequent mutating method calls. A similar, but more specific, optimization is realized by the “use standard charset” action type. This action type replaces references to Java’s standard charsets with the canonical named constants, and serves as an example of integrating domain logic into the serialization process. To further facilitate extensibility, such custom adapters and action types can be provided as plugins.
Listing 8: Plain code serialization of a Monkey, mixing structure- and trace-based serialization
@@ 2.4.4 Implementation
ProDJ relies on Spoon [19] for source code analysis, in order to identify the associated types of methods involving serialization within the target Java application. ProDJ uses Z3 [20] to solve the constraint problem of subsubsection 2.2.2 and determine the most idiomatic reconstruction plan for types that are structure-based serializable. The event sequences for trace-based serializable objects are logged in JSON. The instrumentation of methods and their associated types is achieved using ByteBuddy [^3] https://bytebuddy.net/.
@@ 3.2 Subjects
PDFBox 8f23f8791 170,823 1,390 738 183 188
GraphHopper f64c57c2e 96,716 945 590 199 114
BroadLeaf 951216e942 198,602 2,933 1,883 508 444
Gephi 82684f927 130,705 1,057 1,721 140 225
@@
Table 2: Projects used in the evaluation of plain code serialization: the exact version (SHA) we use for our experiments, the number of lines of Java code (LOC), the number of serialization points we consider (#SPs), the number of types related to the serialization points (#Types), the number of serialization triggered in production by our workloads (#SPs Prod), and the number of types that we target for plain code serialization at runtime (#Types Prod).
@@
For each project in Table 2, column $2$ links to the exact version (SHA) we use for our evaluation, and column $3$ presents the number of lines of Java source code within each project, calculated with tokei. Columns $4$ and $5$ give the number of methods instrumented by ProDJ within the project (#SPs for ‘serialization point’), and consequently, the number of associated types (#Types) that need to be instrumented at this point. For example, ProDJ adds serialization points in $2,933$ methods within BroadLeaf, involving $2,223$ types. For our evaluation, we select methods within each project that are public, non-abstract, non-static, non-deprecated, contain more than one statement, and are declared in a public, non-anonymous, non-local, and non-deprecated class. Columns $6$ and $7$ of Table 2 provide the same metrics, but computed after runtime observations, i.e., the number of serialization points and types that are covered at runtime, which we discuss in the next subsection.